OpenStudy (anonymous):

Divide. 12 divided by 1/6 =

OpenStudy (anonymous):

Here is the next one: 4 divided by 1/9 =

OpenStudy (anonymous):

4 /1 / 9 = 36

OpenStudy (anonymous):

= 12 x 6/1 = 72

OpenStudy (anonymous):

12 divided by 1/6 is 72. 4 divided by 1/9 is 36.

OpenStudy (anonymous):

Thank you so much everyone! I am closing this question because I got my answers from all you wonderful people! Thanks again! ~MA ( MusicalAnimal)